OpenStudy (anonymous):

Two polynomials P and D are given. Use either synthetic or long division to divide P(x) by D(x), and express P in the form P(x) = D(x) · Q(x) + R(x).

OpenStudy (anonymous):

P(x) = x4 − x3 + 9x + 3, D(x) = x2 + 8 P(x) = (x2 + 8) ( ) + ( )

OpenStudy (anonymous):

Who can solve it?

OpenStudy (goten77):

technically theirs infinite solutions... but when doing the long divisions part that restricts it......sadly i forgot how to do this atm.. but ill think for a minute

OpenStudy (callisto):

x^2 ---------------------- x^2 + 8 ) x^4 − x^3 + 9x + 3 x^4 +8x^2 ---------------------- - x^2 - 8x^2 Can you continue...?